jjprincess:

What is the sum of (–2.1x 3.7) and (5 4.9x)?

jjprincess:

someone help please

jjprincess:

someone help please

Olivia16:

When you use sum you add -2.1x + 3.7 + 5 + 4.9x. Now you add the variables and the constants -2.1x + 4.9x = 2.8x 3.7 + 5 = 8.7

jjprincess:

not understanding

Olivia16:

Ok so look at your equation the SUM of, so your adding..-2.1x + 4.9x = 2.8x 3.7 + 5 = 8.7 so it would be 2.8x+8.7

jjprincess:

thank you

jjprincess:

that was very helpful

Olivia16:

Anytime:)

Olivia16:

Anytime:)